1). What solution (anion) can you add to each of the following cation mixtures to precipitate one cation while keeping the other cation in solution? (refer the solubility table)

a). Ba2+ (aq) and Mg2+ (aq)

b). Cu2+ (aq) and Hg22+ (aq)

c). NH4+ (aq) and Ca2+ (aq)

d). Fe2+ (aq) and Pb2+ (aq)

e). Ag+ (aq) and Zn2+ (aq)

Can someone explain to me, in simple terms, how these problems are solved?
Answer #1

a). Ba2+   (aq) and Mg2+ (aq)

Ba + K2CrO4 --> BaCrO4(s) + 2K+

b). Cu2+ (aq) and Hg22+ (aq)

addition of HCl will preciptiate Hg2+2 and not Cu+2

Hg2+2 +2 Cl- = Hg2Cl2(s)

c). NH4+ (aq) and Ca2+ (aq)

addition of Na3PO4 will preicpitate Ca2+ + PO4-3 = Ca3(PO4)3(s)

d). Fe2+  (aq) and Pb2+ (aq)

addition of HCl will form

Pb+2 + 2Cl- = PbCl2(s)

e). Ag+ (aq) and Zn2+ (aq)

Ag+ precipitates with HCl to form

Ag+ + Cl- = AgCl(s)